OpenStudy (lxelle):

given that y(2-x)=3, show that 3 d2y/dx2 - 2y dy/dx = 0

OpenStudy (lxelle):

@ganeshie8 Not able to find? :(

ganeshie8 (ganeshie8):

y(2-x) = 3 differentiate both sides with respect to x and get y'(2-x) + y(-1) = 0

ganeshie8 (ganeshie8):

replace 2-x by 3/y : y'*3/y - y = 0 multiply y through out 3y' - y^2 = 0

ganeshie8 (ganeshie8):

differentiate through out again and you're done!

OpenStudy (lxelle):

THANKS!!!! I have another question! Given that y=(1+4x)e^-2x , show that d2y/dx2 + 4 dy/dx + 4y = 0

OpenStudy (lxelle):

@ganeshie8

ganeshie8 (ganeshie8):

y = (1+4x)e^-2 y' = ?

OpenStudy (samigupta8):

U have to follow the same method of differentiation in this problem as was described by @ganeshie8

OpenStudy (anonymous):

\[y e{2x}=1+4x\] diff. twice w.r.t. x \[y~e ^{2x}2+e ^{2x}\frac{ dy }{ dx }=4\] \[2~y~e ^{2x}2+2~e ^{2x}\frac{ dy }{ dx }+e ^{2x}\frac{ d^2y }{ dx^2 }+2~e ^{2x}\frac{ dy }{ dx }=0\] divide by \[e{2x}\] \[4y+4\frac{ dy }{ dx }+\frac{ d^2y }{ dx^2 }=0\]